### v1

# Black Hole Superradiance of Interacting Multi-Field

Abstract
We investigate black hole superradiance evolution of the interacting multiple fields. We consider a model of two scalar fields interacting with a cubic coupling, and study the superradiant evolution of the cloud. We demonstrate that superradiance is typically suppressed when the superradiant field couples to another field, even with a very weak coupling strength. This implies that the constraints on dark particles derived from single-field analyses can be revised in the presence of interactions. Moreover, we find that the multi-field superradiant evolution and its corresponding observational signatures can be different across parameter spaces, which makes black hole superradiance an even more powerful probe of the dark sector in particle physics.
I Introduction
Bosonic fields around a rapidly rotating black hole manifest an instability known as black hole superradiance [1, 2, 3, 4, 5]. In particular, ultralight bosons can extract angular momentum and energy from an astrophysical black hole, forming a massive cloud around it. With recent advances in gravitational wave detection and black hole imaging, black hole superradiance has emerged as a powerful probe of the dark sector in particle physics [6, 7]. Extensive studies have explored the observational consequences of black hole superradiance, including black hole spin down [8], gravitational wave emission from the clouds [9, 10, 11], and dynamical signatures of clouds in black hole binaries [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33], leading to stringent constraints on ultralight bosonic fields [34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44].
While the dark sector may consist of multiple particle species with non-trivial interactions, most studies assume the presence of only a single superradiant field. Although some works have considered interactions between the superradiant field and other fields, their focus has largely been on phenomena, such as photon polarization by the cloud [45, 46, 47] and annihilation of the cloud into photons [48], fermions [49] and scalars [50], with the assumption of a preformed superradiant cloud. Not only has the superradiance of the coupled fields (or generally, interactions between the black hole and bound states of the coupled fields) been overlooked, but its consequential effects on the superradiance of the primary field also lack investigation. Therefore, a systematic investigation of superradiant evolution of the interacting multiple fields is essential and necessary.
In this work, we investigate black hole superradiance in the presence of multiple fields. We consider a model of two scalar fields interacting with a cubic coupling, and study the superradiance evolution of the cloud. We demonstrate that superradiance is typically suppressed if the superradiant field couples to another field even with a very weak coupling strength. This implies that the constraints on dark particles derived from single-field analyses can be significantly revised in the presence of interactions. Moreover, the superradiant evolution in different regions of the parameter space are distinct. This makes black-hole superradiance an even more powerful probe of the dark sector in particle physics.
The rest of this paper is organized as follows. We start with the superradiance evolution of two non-interacting fields in Sec. II. Then we shall consider a cubic interaction, and shall discuss its effects on the superradiant growth rate in Sec. III and the superradiant evolution in Sec. IV. Finally, we shall discussion the implications observations in Sec. V. Sec. VI devotes to conclusion and discussion. We will take (,+,+,+)(-,+,+,+) metric convention and set $\hbar=c=1$.
II Superradiance of free multi-fields
In this section, we investigate the superradiance process of non-interacting multi-fields. For the purpose of demonstration, we shall start with a model of two massive scalar fields in the Kerr background, the Lagrangian of which is given by
${\cal L}=-\frac{1}{2}g^{ab}\partial_{a}\psi\,\partial_{b}\psi-\frac{1}{2}\mu^{2}\psi^{2}-\frac{1}{2}g^{ab}\partial_{a}\varphi\,\partial_{b}\varphi-\frac{1}{2}\nu^{2}\varphi^{2}+{\cal L}_{\rm int}\,$, (1)
where $g_{ab}$ is the Kerr metric, and $\mu$ and $\nu$ denote the mass of fields $\psi$ and $\varphi$ respectively. Here we have also included ${\cal L}_{\rm int}$, denoting the interactions between the two fields. We shall take ${\cal L}_{\rm int}=0$ in this section, and will consider the interactions in the later sections. For convenience, we define $\alpha\equiv GM\mu$, and the mass ratio $q=\nu/\mu$.
II.1 Eigenstates of free fields
Without direct interactions, the fields only talk to the background, and could be unstable on the Kerr background. The instability manifests in the eigenfrequencies of the fields. Taking $\psi$ for instance, the Klein-Gordon equation satisfied by $\psi$ is separable with the ansatz
$\psi\left(x^{\mu}\right)=e^{-i\omega t}R_{\ell m}(r)S_{\ell m}(\theta)e^{im\varphi}\,$, (2)
where $x^{\mu}=(t,r,\theta,\varphi)$ are the Boyer-Lindquist coordinates, $R_{\ell m}(r)$ is the radial function, and $S_{\ell m}(\theta)$ turns out to be the spheroidal harmonics. The boundary conditions at the black hole horizon and at spatial infinity single out a set of eigenfrequencies. For bound states, i.e., states vanishes at spatial infinity, the eigenfrequencies $\omega$ are shown to be a set of discrete complex number $\omega_{n\ell m}$, labelled by three quantum numbers $n$, $\ell$ and $m$. These eigenfrequencies can be obtained numerically. For instance, Fig. 1 shows the eigenfrequencies of some bound states obtained with the continued fraction method [51]. For $\alpha\ll 1$, we have [14]
${\rm Re}\,\omega_{n\ell m}=\mu\left(1-\frac{\alpha^{2}}{2n^{2}}-\frac{\alpha^{4}}{8n^{4}}+\frac{(2\ell-3n+1)\alpha^{4}}{n^{4}(\ell+1/2)}\right$. (3)
$\left.+\frac{2a_{*}m\alpha^{5}}{n^{3}\ell(\ell+1/2)(\ell+1)}+{\cal O}(\alpha^{6})\right)\,$,
and [4, 52, 53]
${\rm Im}\,\omega_{n\ell m}\approx 2\tilde{r}_{+}(m\Omega_{H}-\omega_{n\ell m})\alpha^{4\ell+5}C_{n\ell m}$ (4)
with
$C_{n\ell m}=\frac{2^{4\ell+2}(2\ell+n+1)!}{(n+\ell+1)^{2\ell+4}n!}\left(\frac{\ell!}{(2\ell)!(2\ell+1)!}\right)^{2}$ (5)
$\times\prod^{\ell}_{j=1}\left(j^{2}(1-a^{2})+(ma-2\tilde{r}_{+}\alpha)^{2}\right)\,$,
where $a$ is the black hole dimensionless spin, $\tilde{r}_{+}\equiv 1+\sqrt{1-a^{2}}$, and $\Omega_{H}\equiv a/2M\tilde{r}_{+}$ is the angular velocity of the outer horizon. [^1] Also see Refs. [54, 55] for the improved analytical superradiance solutions. Therefore, when $m\Omega_{H}>\omega_{n\ell m}$, the bound state experiences the superradiant growth with a rate of $\Gamma_{n\ell m}\equiv{\rm Im}\,\omega_{n\ell m}$.
Figure 1: Eigenfrequencies of bound states with $n\ell m$ being 211211 (in black) and 322322 (in red). The upper and lower plots show the real and imaginary parts of the eigenfrequencies respectively. The solid lines are given by Eqs. (3) and (4), while the dashed lines show the resultes obtained from numerical calculation.
While the wavefunction of the bound states can also be obtained numerically [56], for $\alpha\ll 1$, it is convenient to work with the wavefunction in the non-relativistic limit. Taking the ansatz
$\psi=\frac{1}{\sqrt{2\mu}}\left({\bar{\psi}}e^{-i\mu t}+c.c.\right)\,$, (6)
where $\bar{\psi}$} is a complex field that varies on a time scale much longer than $\mu^{-1}$, the Klein-Gordon equation reduces to a Schrödinger-like equation,
$\left[i\partial_{t}+\frac{\nabla^{2}}{2\mu}+\frac{\alpha}{r}+{\cal O}(\alpha^{2})\right]{\bar{\psi}}=0\,$, (7)
and the eigenstates of $\bar{\psi}$} is given by
${\bar{\psi}}_{n\ell m}=\sqrt{N_{n\ell m}}\,u_{n\ell m}({\bf r})e^{-i\delta_{n\ell m}t}\,$, (8)
where $\delta_{n\ell m}\equiv\omega_{n\ell m}-\mu$, $N_{n\ell m}$ is the occupation number of particle in the eigenstate, and $u_{n\ell m}({\bf r})$ is normalized to 1. To the leading order in $\alpha$, $u_{n\ell m}({\bf r})$ is given by the normalized hydrogenic eigen-wavefunction with a Bohr radius of $r_{B}\equiv GM/\alpha^{2}$.
The Klein-Gordon equation of $\psi$ also allows unbound states, which satisfy the out-going boundary condition at spatial infinity. The unbound states are continuous in spectrum, with the eigenfunctions, in the Newtonian limit, given by the stationary Coulomb waves [57]
$\psi_{n\ell m}\propto u_{k\ell m}({\bf r})e^{-i\omega_{k\ell m}t}\,.$ (9)
The above discussion can be easily extend to the $\varphi$ field by replacing $\alpha$ with $q\alpha$.
Figure 2: Parameter space of two non-interacting fields. Without loss of generality, we assume $\psi_{211}$ to be the fastest growing mode, which excludes the white regime. Representative examples of the evolution in different colored regions are shown in Fig. 3. In particular, we find that superradiant growth of $\psi_{211}$ might be affect by $\varphi$, as which may accelerate black hole spin, causing $\psi_{211}$ depletes prematurely.
Figure 3: Presentative examples of non-interacting two-field superradiant evolution. The upper panel shows the evolution of occupation numbers in the region of corresponding color in Fig. 2, while the lower panel shows the evolution of black hole spin.
II.2 Superradiance Evolution
The superradiant growth of the fields back reacts on the background geometry, leading to black hole spin down. Under the adiabatic approximation, the spin and mass of the black hole evolve as
$\dot{a}=-\sum_{i}m_{i}\gamma_{i}\epsilon_{i}$ (10)
$\frac{\dot{M}}{\mu}=-GM^{2}\sum_{i}\gamma_{i}\epsilon_{i}\,$, (11)
where we have defined $\gamma_{i}=\Gamma_{i}/\mu$ and $\epsilon_{i}=N_{i}/GM^{2}$, the over-dot denotes derivative with respect to the dimensionless time $\tilde{t}\equiv\mu t$, and the summation over all bound states of both $\psi$ and $\varphi$.
In the case of single filed $\psi$, $\psi_{211}$ typically grows most rapidly with
$\gamma_{211}=\frac{1}{24}\alpha^{8}\left[a-2\alpha\left(1+\sqrt{1-a^{2}}\right)\right]\,$, (12)
and first dominates the black hole spin evolution. It continues growing until the black hole spin drops below $a_{\rm crit}=4\alpha/(1+4\alpha^{2})$, namely the superradiant condition is no longer satisfied. At this time, $\psi_{211}$ saturates with an occupation number of $\epsilon_{\rm max}=(a_{0}-a_{\rm crit})\sim 1$.
In the case of two fields, the superradiance evolution depends on the parameters. In Fig. 2, we show the parameter space of different types of evolution, assuming $\psi_{211}$ to be the fastest growing mode, [^2] The case that $\varphi_{211}$ grows faster can be obtained by replacing $q$ with $1/q$. while a representative example of each type is shown in Fig. 3. These examples are obtained by solving Eq. (10) together with $\dot{\epsilon}_{i}=\gamma_{i}\epsilon_{i}$, assuming the initial conditions of $\epsilon_{i}=10^{-10}$ when the black hole spin $a_{0}=0.99$. For simplicity, we treated the black hole mass (and hence $\alpha$) to be constant, given the fact that the black hole mass typically varies by a small amount during superradiant evolution. We find that, in region I and region IV, $\psi_{211}$ finally depletes after reaching the maximum occupation number, due to the development of $\varphi$ field.
III Cubic Interactions
In this section, we investigate the superradiant growth of interacting scalar fields. We shall consider cubic interactions, which might take effects first comparing to higher dimensional interactions. We shall focus on the interactions between the fields instead of self-interactions, as which have been investigated in Ref. [8]. This leaves two relevant operators: $\psi^{2}\,\varphi$ and $\psi\,\varphi^{2}$. As the growth rate is extremely sensitivity to the mass of the fields, one of the fields usually grow much faster than another field, even if the masses of the two field are comparable. Without loss of generality, let $\psi$ be the one that grows faster, in which case $\psi^{2}\,\varphi$ is expected to dominate over $\psi\,\varphi^{2}$. For these reasons, we shall neglect $\psi\,\varphi^{2}$, and consider Lagrangian (1) with ${\cal L}_{\rm int}=\lambda\psi^{2}\,\varphi$, where $\lambda$ is the coupling constant.
The field equations are
$\left(\Box-\mu^{2}\right)\psi=-2\lambda\psi\varphi\,$, (13)
$\left(\Box-\nu^{2}\right)\varphi=-\lambda\psi^{2}\,$, (14)
where $\Box$ denotes the D’Alembert operator in the Kerr background. We shall further assume the two fields are weakly coupled.
Assuming both fields are initially in vacuum states with small quantum fluctuations, the superradiance process can be investigated perturbatively: At the very beginning, interactions between the field is suppressed due to the low occupation numbers, and the superradiance process is the same as it in the case of free field. As superradiance continues, the occupation number in the fastest growing mode, i.e., ${\bar{\psi}}_{211}$, increases, resulting in a notable interaction between the two fields. Depending on the parameters, the interaction may lead to different processes. The effects on the superradiant growth of such processes can be investigated by their corrections on the eigenfrequncies, which will be discussed as follows.
We shall take the non-relativistic ansatz (6) as in the case of free field, and work in the weak field limit. Then Eqs. (13) and (14) can be written as
$\left(i\partial_{t}+\hat{H}\right)\left({\bar{\psi}}e^{-i\mu t}+{\bar{\psi}}^{*}e^{i\mu t}\right)=-2\bar{\lambda}\left({\bar{\psi}}e^{-i\mu t}+{\bar{\psi}}^{*}e^{i\mu t}\right)\varphi$ (15)
$\left(\partial_{t}^{2}-\hat{M}-\nu^{2}\right)\varphi=\bar{\lambda}\left({\bar{\psi}}^{2}e^{-2i\mu t}+2{\bar{\psi}}{\bar{\psi}}^{*}+{{\bar{\psi}}}^{*2}e^{2i\mu t}\right)\,$,
where $\bar{\lambda}\equiv\lambda/2\mu$ is the reduced coupling strength, and $\hat{M}\approx-\nabla^{2}-(2q\alpha\nu)/r$ to the first order in $GM/r$. Although working in the weak field limit, it is useful to consider $\hat{H}$ and $\hat{M}$ as non-Hermintian operators in order to take into account the energy exchanges between the fields and the black hole that do not manifest in the weak field limit.
To compute the corrections on the eigenfrequency, we take one of the $\bar{\psi}$} in the source terms as background, i.e., ${\bar{\psi}}_{0}=\sqrt{N_{0}}u_{0}(\mathbf{r})e^{-i\delta_{0}t}$. While ${\bar{\psi}}_{0}$ is in the ${\bar{\psi}}_{211}$ mode which should in principle grow with time, we shall consider ${\bar{\psi}}_{0}$ to be stationary in our perturbative calculation, and take $\delta_{0}$ to be the real part of $\omega_{211}-\mu$. It is convenient to use Feynman-like diagrams to represent the channels involved in Eqs. (LABEL:eq:eomm). In particular, $\bar{\psi}$} or $\bar{\varphi}$} on the r.h.s. of Eqs. (LABEL:eq:eomm) corresponds to a leg on the l.h.s. of the vertex, and vice versa, while $\bar{\psi}$$^{*}$ or $\bar{\varphi}$$^{*}$ on the r.h.s. of Eqs. (LABEL:eq:eomm) corresponds to a leg on the r.h.s. of the vertex, and vice versa. Here $\bar{\varphi}$} is defined as $\varphi={\bar{\varphi}}+{\bar{\varphi}}^{*}$. Focusing on interaction corrections on the fastest growing mode ${\bar{\psi}}_{211}$, Eqs. (LABEL:eq:eomm) involve three possible channels, which are shown in Fig. 4.
$\psi_{211}$$\psi_{211}$$\psi_{211}$$\psi_{211}$
$\psi_{211}$$\psi_{211}$$\psi_{211}$$\psi_{211}$
$\psi_{211}$$\psi_{211}$$\psi_{211}$$\psi_{\infty}$
Figure 4: Possible channels involved in Eqs. (LABEL:eq:eomm). From left to right, the diagrams correspond to the s-channel, the t/u channel and the $3{\bar{\psi}}$ channel. As we are considering self-interaction correction, we do not distinguish t and u channel in the diagram.
Figure 5: Numerical results of integrals (24), (26), and (37). In practice, we have to impose a cut-off $n_{\rm max}$ for the summation in Eq. (23). Given the left plot, we expect the summation converges well when $n_{\rm max}>8$ for $q<5$.
Taking the s-channel for instance, we have
$\left(i\partial_{t}+\hat{H}\right){\bar{\psi}}=-2\bar{\lambda}{\bar{\psi}}_{0}^{*}{\bar{\varphi}}e^{2i\mu t}$ (16)
$\left(\partial_{t}^{2}-\hat{M}-\nu^{2}\right){\bar{\varphi}}=\bar{\lambda}{\bar{\psi}}{\bar{\psi}}_{0}e^{-2i\mu t}\,$,
to solve which, we can take the ansatz
${\bar{\psi}}(t,\mathbf{r})$ $=\Psi(\mathbf{r})e^{-i(\delta_{0}+\delta\omega_{211})t}$ (17)
$\bar{\varphi}(t,\mathbf{r})$ $=\Phi(\mathbf{r})e^{-i\sigma t}\,.$
To the zeroth order of $\bar{\lambda}$, we have $\Psi=u_{0}$ and $\Phi=0$. To the first order of $\bar{\lambda}$, we have $\sigma=2\omega_{0}+\delta\omega_{211}$ and
$\left(\epsilon_{0}+\delta\omega_{211}+\hat{H}\right)\Psi=-2\bar{\lambda}\sqrt{N_{0}}u_{0}^{*}\Phi$ (18)
$\left(-\sigma^{2}-\hat{M}-\nu^{2}\right)\Phi=\bar{\lambda}\sqrt{N_{0}}u_{0}\Psi\,.$
In order to solve $\Phi$, we further wirte
$\Phi=\sum_{n,\ell,m}^{\infty}b_{n\ell m}v_{n\ell m}+\sum_{\ell,m}\int_{0}^{\infty}b_{\ell m}(k)v_{k\ell m}\mathrm{d}k\,$, (19)
where $v_{n\ell m}$ and $v_{k\ell m}$ are the eigenfunctions satisfying
$\hat{M}v_{n\ell m}=\kappa_{n}^{2}v_{n\ell m},\quad\hat{M}v_{k\ell m}=-\kappa^{2}v_{k\ell m}$ (20)
with $\kappa_{n}\equiv\sqrt{\nu^{2}-\sigma_{n}^{2}}$ and $\kappa\equiv\sqrt{\sigma^{2}-\nu^{2}}$ being the eigenvalues. Then $\Phi$ in Eqs. (LABEL:eq:sschannel) can be solved with the orthogonal conditions. Specifically, integral over the angular part indicates the nontrivial contributions only come from the modes with $\ell=m=2$, and their corresponding coefficients are
$b_{n}$ $=\frac{\bar{\lambda}\sqrt{N_{0}}}{-(2\omega_{0})^{2}-\kappa_{n}^{2}+\nu^{2}}\int u_{0}^{2}v_{n22}^{*}\mathrm{d}^{3}\mathbf{r}$ (21)
$b(k)$ $=\frac{1}{2\pi}\frac{\bar{\lambda}\sqrt{N_{0}}}{-(2\omega_{0})^{2}+\kappa^{2}+\nu^{2}}\int u_{0}^{2}v_{k22}^{*}\mathrm{d}\,,^{3}\mathbf{r}$
where we have omitted the subscriptions $\ell m$} for short. Then by multiplying $u_{0}^{*}$ on the equation of $\Psi$ and integrating over $\mathrm{d}^{3}\mathbf{r}$ on both sides, we find $\Psi=u_{0}$ to the leading order of $\bar{\lambda}$. Also, the frequency correction from the s-channel can be written as $\delta\omega_{211}^{s}=\delta\omega_{211}^{sB}+\delta\omega_{211}^{sE}$ with
$\delta\omega_{211}^{sB}=2\bar{\lambda}^{2}N_{0}\sum_{n\geq 3}\frac{\left|\int u_{0}^{2}v_{n22}^{*}\mathrm{d}^{3}\mathbf{r}\right|^{2}}{k_{0}^{2}+\kappa_{n}^{2}}$ (22)
$\delta\omega_{211}^{sE}=2\bar{\lambda}^{2}N_{0}\int_{0}^{\infty}\frac{\mathrm{d}k}{2\pi}\,\frac{\left|\int u_{0}^{2}v_{k22}^{*}\mathrm{d}^{3}\mathbf{r}\right|^{2}}{k_{0}^{2}-\kappa^{2}(k)}\,$,
representing the corrections from bound states and unbound states of $\varphi$ respectively. Here, we have defined $k_{0}^{2}\equiv 4\omega_{0}^{2}-\nu^{2}$, and $k\equiv\mathrm{Re}\,\kappa$.
We are interested in the interaction corrections on the growth rate, which can be inferred from the imaginary part of $\delta\omega_{211}^{s}$. For the corrections from bound states, we find
${\rm Im}\delta\omega_{211}^{sB}\simeq\alpha^{3}\bar{\lambda}^{2}N_{0}\frac{4q^{2}\mu}{(q^{2}-4)^{2}}\sum_{n\geq 3}\frac{{\rm Im}\sigma_{n}}{\nu}{\rm I}_{n}^{B}(q)\,$, (23)
which indicates that the superradiant growth rate of ${\bar{\psi}}_{211}$ is suppressed if it interacts with a decaying ${\bar{\varphi}}_{n22}$, as ${\bar{\varphi}}_{n22}$ may lose energy to the black hole in the s-channel, and vice versa. Here
${\rm I}_{n}^{B}\equiv\frac{1}{\mu^{3}\alpha^{3}}\left|\int u_{0}^{2}v_{n22}^{*}\mathrm{d}^{3}\mathbf{r}\right|^{2}\,$, (24)
is a dimensionless factor depending only on the mass ratio $q$, with its values shown in Fig. 5. For the corrections from unbound states, we find that $\delta\omega_{211}^{sE}$ has a non-zero imaginary part only if $k_{0}$ is real due to the optical theorem. Specifically, one has
${\rm Im}\delta\omega_{211}^{sE}\simeq\left\{\begin{array}[]{ll}-\bar{\lambda}^{2}N_{0}\frac{\mu\alpha^{2}}{2\sqrt{4-q^{2}}}{\rm I}^{E}_{k_{0}}(q,\alpha),&2\omega_{0}>\nu\\ 0,&2\omega_{0}<\nu\end{array}\right$. (25)
which represents the superradiant growth of ${\bar{\psi}}_{211}$ is suppressed by their annihilation into radiation of $\varphi$, a process happening only if $2\omega_{0}\geq\nu$. Here
${\rm I}^{E}_{k_{0}}\equiv\frac{1}{\mu^{2}\alpha^{2}}\left|\int u_{0}^{2}v_{k_{0}22}^{*}\mathrm{d}^{3}\mathbf{r}\right|^{2}$ (26)
is a dimensionless factor, with its values shown in Fig. 5.
One should notice that the above calculation is valid only if the interaction terms are perturbative, in other words, when $|\mathrm{Re}\,\delta\omega|\ll|\delta_{0}|$. Since $\mathrm{Re}\,\delta\omega$ is dominated by $\delta\omega^{sB}$, we can define a threshold particle number
$N_{\rm th}\equiv\left|\frac{q^{2}-4}{2\bar{\lambda}^{2}\alpha\sum_{n}I_{n}^{B}}\right|$ (27)
beyond which the interaction effects cannot be treated perturbatively. In addition, there are t and u channels, which can be investigated in a similar way. The frequency corrections, however, happen to be real, and therefore will not be discussed in details.
Besides radiating $\varphi$, the superradiant mode could also lose energy to spacial infinity by radiating $\psi$ via a process depicted by the right diagram in Fig. 4. This process, hereinafter referred to as the $3{\bar{\psi}}$-process, is described by
$\left(\Box-\mu^{2}\right)\psi=-2\lambda\sqrt{\frac{N_{0}}{2\mu}}u_{0}e^{-i\omega_{0}t}{\bar{\varphi}}+c.c.\,$, (28)
$\left(\Box-\nu^{2}\right){\bar{\varphi}}=-\bar{\lambda}N_{0}u_{0}^{2}e^{-2i\omega_{0}t}\,$, (29)
which are obtained from Eqs. (13) and (14), with $\psi$ on the r.h.s. being the background ${\bar{\psi}}_{0}$. The effects of $3{\bar{\psi}}$-process on the growth rate of ${\bar{\psi}}_{211}$ can be investigated by estimating the radiation power of $\psi$.
Figure 6: Parameter space of interaction corrections. From left to right, the plots show the regions of nontrivial $\delta\omega_{211}^{sE}$ (in red), $\delta\omega_{211}^{sB}$ (in blue), and $\Gamma_{R}$ (in green). Again, we exclude the white regime as we assume ${\bar{\psi}}_{211}$ to be the fastest growing mode. Moreover, we further distinguish the regions that ${\rm Im}\delta\omega_{211}^{sB}>0$ (in dark blue), ${\rm Im}\delta\omega_{211}^{sB}<0$ (in light blue), and ${\rm Im}\delta\omega_{211}^{sB}$ changes from positive to negative as the black hole spin evolves from $\sim 1$ to $a_{crit}$ (in blue).
Again, we express $\varphi$ with its free eigenfunctions,
${\bar{\varphi}}=\left(\sum_{n\geq 3}c_{n}\,v_{n22}+\int{\rm d}k\,c(k)\,v_{k22}\right)e^{-2i\omega_{0}t}$ (30)
and with Eq. (29) we get
$c_{n}=-\frac{\bar{\lambda}N_{0}}{k_{0}^{2}+k_{n}^{2}}\int u_{0}^{2}v_{n22}^{*}\,\mathrm{d}^{3}\mathbf{r}$ (31)
$c(k)=\frac{1}{2\pi}\frac{\bar{\lambda}N_{0}}{k^{2}-k_{0}^{2}}\int u_{0}^{2}v_{k22}^{*}\,\mathrm{d}^{3}\mathbf{r}\,$,
where we have considered the fact that only modes with $\ell=m=2$ contribute. Substituting Eq. (30) into Eq. (28) leads to
$\left(\Box-\mu^{2}\right)\psi=f(\mathbf{r})e^{-i\omega_{R}t}+c.c.$ (32)
where $\omega_{R}=3\omega_{0}$ is the driving frequency, and
$f(\mathbf{r})\equiv-2\lambda\sqrt{\frac{N_{0}}{2\mu}}u_{0}\left(\sum_{n}c_{n}\,v_{n22}+\int{\rm d}k\,c(k)\,v_{k22}\right)\,.$ (33)
Then the radiation power of $\psi$ is given by
[…]
where $\mathrm{d}\Omega_{k}$ is the differential solid angle pointing the direction $(\theta_{k},\phi_{k})$, and
$\tilde{f}(\mathbf{k})=4\pi\sum_{\ell m}Y_{\ell m}(\theta_{k},\phi_{k})\int\mathrm{d}^{3}\mathbf{r}\,(-i)^{\ell}f(\mathbf{r})\frac{u^{*}_{k\ell m}}{2k}\,.$ (35)
The angular integral in Eq. (35) indicates that only $u^{*}_{k\ell m}$ with $\ell=m=3$ contributes, and eventually the radiation power can be written as
$P=\frac{4\omega_{R}\mu^{2}}{\sqrt{\omega_{R}^{2}-\mu^{2}}}N_{0}^{3}\bar{\lambda}^{4}\alpha^{5}{\rm I}^{R}(\alpha,q)$ (36)
where
${\rm I}^{R}(\alpha,q)\equiv\frac{1}{\alpha^{5}\mu}\left|\sum_{n}\frac{\int u_{0}^{2}v_{n22}^{*}\mathrm{d}^{3}\mathbf{r}}{-k_{n}^{2}-k_{0}^{2}}\int u_{0}v_{n22}u_{k\ell m}^{*}\mathrm{d}^{3}\mathbf{r}\right$. (37)
$\left.+\int\frac{\mathrm{d}k}{2\pi}\frac{\int u_{0}^{2}v_{k22}^{*}\mathrm{d}^{3}\mathbf{r}}{k^{2}-k_{0}^{2}}\int u_{0}v_{k22}u_{k\ell m}^{*}\mathrm{d}^{3}\mathbf{r}\right|^{2}$
is a dimensionless quantity including all summations and integrals. The numerical value of some ${\rm I}^{R}(\alpha,q)$ is shown in Fig. 5. Because of energy conservation, the radiation turns on only if $3\omega_{0}\geq\mu$ which is always satisfied providing $\psi$ is a superradiant field. Given the energy of each radiated particle is of $\omega_{R}=3\omega_{0}$, we can further define a decay rate given the radiation power,
$\Gamma_{R}\equiv\frac{P}{\omega_{R}N_{0}}\simeq\sqrt{2}\mu N_{0}^{2}\bar{\lambda}^{4}\alpha^{5}{\rm I}^{R}(\alpha,q)\,.$ (38)
The parameter space for the processes discussed above are summarized in Fig. 6.
Figure 7: Parameter space of the dominating interaction corrections. We show the regions that the term of $\gamma_{E}$ (in red), the $\gamma_{B}$ (in blue) and $\gamma_{R}$ (in green) might dominate the superradiant evolution. We further exclude the region that $\epsilon_{eq}>\epsilon_{th}\equiv N_{th}/GM^{2}$, cf. Fig. 6, where the perturbative treatment may break down. Moreover, we also show the contours of $\tilde{\lambda}_{c}$ in each color regions. We expect that the interaction correction can interrupte the superradiant growth of $\psi_{211}$ significantly, and lead to the quasi-equilibrium stage if $\tilde{\lambda}>\tilde{\lambda}_{c}$. Given the value of $\tilde{\lambda}_{c}$, we find that the superradiant growth of ${\bar{\psi}}_{211}$ can be easily affected by $\varphi$ even with very weak coupling, i.e., $\lambda/\mu\ll 1$.
IV Superradiance Evolution of interacting fields
Figure 8: The three stages of interacting two-field superradiant evolution. The left and right plots show the evolution of the dimensionless occupation number $\epsilon_{211}$ and the black hole spin respectively, where the vertical grids denotes $\tau_{sr}$ defined in Eq. (44), while the horizontal grids denote $\epsilon_{eq}$ and $a_{crit}$. The gray oblique line in the left plot are given by Eq. (47), showing the asymptotical behavior.
In this section, we discuss superradiance evolution of the ${\bar{\psi}}_{211}$ mode in the presence of the interaction. We shall neglect the effects from the superradiant growth of the other modes, and will justify this treatment later. Given the possible processes discussed in Sec. III, the superradiance evolution can be understood by considering the evolution of the occupation number in ${\bar{\psi}}_{211}$
$\dot{\epsilon}_{211}=\gamma_{211}\epsilon_{211}-\left(\gamma_{B}+\gamma_{E}\right)\epsilon_{211}^{2}-\gamma_{R}\epsilon_{211}^{3}$ (39)
and its effects on the black hole spin
$\dot{a}=-\gamma_{211}\epsilon_{211}+\gamma_{B}\epsilon_{211}^{2}$ (40)
where the over-dot denotes the derivative with respect to $\tilde{t}=\mu t$, and we have defined
$\gamma_{B}=-\frac{2{\rm Im}\delta\omega^{sB}_{211}}{\epsilon_{211}\,\mu},\quad\gamma_{E}=-\frac{2{\rm Im}\delta\omega^{sE}_{211}}{\epsilon_{211}\,\mu},\quad\text{and}\quad\gamma_{R}=\frac{3\Gamma_{R}}{\epsilon_{211}^{2}\,\mu}\,.$ (41)
In principle, the black hole mass also evolves during the superradiance evolution. Nevertheless, we treat the black hole mass to be constant, since the black hole mass typically varies by a small amount during superradiance. This is especially the case in the presence of the interaction, as the cloud tends to grow to a smaller occupation number. With this treatment, $\alpha$ can also be considered as a constant. We could also take into account gravitational wave radiations by appending $\gamma_{GW}\epsilon_{211}^{2}$ on the r.h.s. of Eq. (39) with $\gamma_{GW}=\alpha^{14}/160$ [58, 59]. While we are interesting in the case with $\epsilon_{eq}$ derived from the interaction corrections is less than 1, the gravitational wave radiations leads to an effective $\epsilon_{eq}\gg 1$ due to the weak nature of gravitational interactions, and therefore is irrelevant for the discussion.
Fig. 8 shows a typical evolution in the presence of interaction, which is obtained by solving Eqs. (39) and (40) numerically. We find that the superrdiant mode typically experiences three stages: exponential growth, quasi-equilibrium, and power-law decay, which can be understood as follows.
Initially, $\epsilon_{211}$ is very small, and its evolution is dominated by the linear term in Eq. (39), in which case the mode will grow exponentially. The exponential growth continues until the $\gamma_{211}$ becomes almost zero due to black hole spin down in which case the cloud saturated, or the interaction terms in Eq. (39) becomes significant. In the former case, the later evolution would be similar to that of free fields. In the later case, the exponential growth could be interrupted by the interaction corrections, and the superradiant cloud researches a quasi-equilibrium stage. By requiring the l.h.s. of Eq. (39) to be zero, one can define the dimensionless occupation number at the quasi-equilibrium stage $\epsilon_{eq}$. We expect the cloud reaches the quasi-equilibrium stage before saturation if $\epsilon_{\rm eq}<\epsilon_{\rm max}$.
Although different processes may take place simultaneously as shown in Fig. 6, typically only one process dominates the correction, which allowing us to simplify the equation by neglecting the sub-dominating correction. In Fig. 7, we show the parameter space of the dominating process, with the contours denoting the critical $\tilde{\lambda}_{c}$ defined as $\epsilon_{eq}(\tilde{\lambda}_{c})=\epsilon_{\rm max}$. Namely, the cloud would reach the quasi-equilibrium stage before it saturates as in the single field case if $\tilde{\lambda}>\tilde{\lambda}_{c}$. Here $\tilde{\lambda}$ relates to $\lambda$ via
$\lambda=\frac{\tilde{\lambda}}{\sqrt{GM^{2}}}\mu\approx\left(\frac{\tilde{\lambda}}{2.7\times 10^{38}}\right)\left(\frac{3M_{\odot}}{M}\right)\mu\,.$ (42)
Given the value of $\tilde{\lambda}_{c}$, we find that the cloud reaches the equilibrium stage with very tiny $\epsilon_{eq}$ even if $\lambda/\mu\ll 1$. Depending on the dominating process, the subsequent evolution may be slightly different, and will be discussed as follows.
Figure 9: Presentative examples of interacting two-field superradiant evolution. The upper panel shows the evolution of occupation numbers in the region of corresponding color in Fig. 6, while the lower panel shows the evolution of black hole spin. The gray dashed line shows the evolution with single ${\bar{\psi}}_{211}$ for comparison.
IV.1 Case A
In region A of Fig. 7, the $\gamma_{E}$ term in Eq. (39) typically dominates over the other correction terms. In this case, we have $\epsilon_{eq}\simeq\gamma_{211}/\gamma_{E}$, and Eq. (39) and Eq. (40) reduce to
$\epsilon_{211}^{\prime}$ $\simeq(\epsilon_{eq}-\epsilon_{211})\epsilon_{211}$ (43)
$a^{\prime}$ $\simeq-\epsilon_{eq}\epsilon_{211}\,$,
where the prime denotes derivative with respect to the rescaled time $\eta=\gamma_{E}\tilde{t}$. The quasi-equilibrium stage can be reached in a time scale of
$\tau_{sr}=\frac{\ln\epsilon_{eq}-\ln\epsilon_{i}}{\gamma_{211}\mu}\,$, (44)
where $\epsilon_{i}$ is the initial value of $\epsilon_{211}$ and should be extremely small. Once reaching the quasi-equilibrium stage, $\epsilon_{211}$ evolves adiabatically following $\epsilon_{eq}(a)$ with the black hole spin satisfying
$a^{\prime}=-\epsilon_{eq}^{2}\,.$ (45)
Without the interruption of other superradiant modes, the quasi-equilibrium stage continues until $\epsilon_{eq}$ approaches to 0 as the black hole spin approaches to the critical spin $a_{crit}$. While Eq. (45) can be solved analytically, we shall not write down the lengthy expression of $a$, but are rather interested in the evolution when $a$ is close to $a_{crit}$, in which case
$\epsilon_{eq}\simeq\frac{\alpha^{8}}{24\gamma_{E}}\left(\frac{1+4\alpha^{2}}{1-4\alpha^{2}}\right)(a-a_{crit})\,.$ (46)
Then Eq. (45) indicates
$a-a_{crit}\simeq\frac{576\gamma_{E}^{2}}{\alpha^{16}}\left(\frac{1-4\alpha^{2}}{1+4\alpha^{2}}\right)^{2}\frac{1}{\eta}\,$, (47)
$\epsilon_{211}\simeq\frac{24\gamma_{E}}{\alpha^{8}}\frac{1-4\alpha^{2}}{1+4\alpha^{2}}\frac{1}{\eta}\,$,
which corresponds the power-law decay as shown in Fig. 8. Note that the other corrections will never dominate the evolution even during the power-law decay.
IV.2 Case B
In region B of Fig. 7, the $\gamma_{R}$ term typically dominates the interaction correction. In this case, we have
$\epsilon^{\prime}_{211}$ $\simeq\epsilon_{\rm eq}^{2}\epsilon_{211}-\epsilon_{211}^{3}$ (48)
$a^{\prime}$ $\simeq-\epsilon_{\rm eq}^{2}\epsilon_{211}.$
with $\epsilon_{eq}\simeq(\gamma_{211}/\gamma_{R})^{1/2}$. While $\epsilon_{211}$ could also reaches $\epsilon_{eq}$ as in case A, the equilibrium may not be always stable. Considering a small deviation $\delta\epsilon$ from $\epsilon_{eq}$ and by perturbing Eq. (48), we have
$\delta\epsilon^{\prime}=-2\epsilon_{\rm eq}^{2}\delta\epsilon+\mathcal{O}(\delta\epsilon^{2}),$ (49)
which means a typical timescale for restoring equilibrium is $\tau_{re}\sim(2\epsilon_{\rm eq}^{2})^{-1}$. On the other hand, the $\epsilon_{eq}$ varies on a time scale of
$\tau_{var}\sim\left(\frac{\mathrm{d}\epsilon_{\rm eq}}{\mathrm{d}a}\epsilon_{\rm eq}^{2}\right)^{-1}\,.$ (50)
Comparing the two time scales $\tau_{var}$ and $\tau_{re}$ leads to a threshold in term of black hole spin
$a_{th}=\frac{\alpha^{8}}{384\gamma_{R}}\frac{1+4\alpha^{2}}{1-4\alpha^{2}}\,$, (51)
below which the quasi-equilibrium stage becomes unstable. If the quasi-equilibrium is stable as $a$ approaches to $a_{crit}$, the cloud decays with $\epsilon_{211}\propto t^{-1}$. Otherwise, $\epsilon^{\prime}_{211}$ is dominated by the $\gamma_{R}\epsilon_{211}^{3}$ term, and the cloud decays with $\epsilon_{211}\propto t^{-1/2}$.
IV.3 Case C
In region C of Fig. 7, the $\gamma_{B}$ term typically first dominates the interaction correction. The evolution is similar to that in case A, except that the black hole spin stops decreasing once reaching the quasi-equilibrium stage. The black hole spin and the occupation number remain almost constant until the $\gamma_{R}$ term becomes significant and takes over the evolution.
IV.4 Cloud Collapse
The cubic interaction can lead to an effective quartic self-interaction of $\psi$, cf. Fig. 4, under which a cloud may collapse as the occupation number grows [8, 7]. The critical occupation number for cloud collapse can be estimated by considering the wavefunction of $\bar{\psi}_{211}$ with a modified radius $R$
$\bar{\psi}_{211}=\frac{\sqrt{N}}{2\sqrt{6}}rR^{-5/2}e^{-r/2R}Y_{11}(\theta,\phi)\,$, (52)
and investigating how $R$ deviates from the Bohr radius $r_{B}$ in the presence of the effective self-interaction. In the non-relativistic limit, the action of $\bar{\psi}$} reduces to
$S_{{\bar{\psi}}}=\int{\rm d}^{3}r{\rm d}t$ $\frac{i}{2}\left(\bar{\psi}^{*}\partial_{t}\bar{\psi}-\bar{\psi}\partial_{t}\bar{\psi}^{*}\right)-\frac{1}{2\mu}|\nabla\bar{\psi}|^{2}-\mu\Phi_{N}|\bar{\psi}|^{2}$ (53)
$+\bar{\lambda}\bar{\psi}^{2}\varphi-\frac{|\nabla\Phi_{N}|^{2}}{8\pi G}-\rho_{\rm BH}\Phi_{N}\,$,
where $\Phi_{N}$ is the gravitational potential satisfying
$\nabla^{2}\Phi_{N}=4\pi G\left(\rho_{\rm BH}+|\bar{\psi}|^{2}\right)\,$, (54)
with $\rho_{\rm BH}=M\delta^{3}({\bf r})$. $\varphi$ is a mediator. Taking into account the contribution of s, t and u channels, we have
$\varphi=$ $\left(\sum_{n\geq 3}c_{n}\,v_{n22}+\int{\rm d}k\,c(k)\,v_{k22}\right)e^{-2i\omega_{0}t}$ (55)
$+\left(\sum_{n\geq 3}d_{n}\,v_{n20}+\int{\rm d}k\,d(k)\,v_{k20}\right)$
$+\left(\sum_{n\geq 1}e_{n}\,v_{n00}+\int{\rm d}k\,e(k)\,v_{k00}\right)+c.c.\,$,
where the coefficients $d_{n},d(k),e_{n},e(k)$ can be obtained similarly as Eq. (31). By inserting Eq. (55) in to action (53) and integrating over space, we can obtain an effective potential of $R$,
[…]
where $\tilde{R}=R/r_{B}$ and $I_{p}$ is a dimensionless factor containing contribution of $\ell m=22,20,00$, with its numerical values of each component shown in Fig. 10. The four terms in Eq. (56) correspond to kinetic energy, black hole gravity, quartic interaction and cloud self-gravity respectively. The extrema of $V_{\rm eff}$ locates at
$\tilde{R}_{\rm m}^{\pm}=\frac{16}{256+711\alpha\epsilon}\left(8\pm\sqrt{64-3I_{p}\alpha\epsilon\tilde{\lambda}^{2}(256+711\alpha\epsilon)}\right).$ (57)
However, when the occupation number reaches
$\epsilon_{\rm collapse}=\frac{8}{711}\sqrt{\frac{237}{I_{p}\tilde{\lambda}^{2}}-256}-\frac{128}{711\alpha}\,$, (58)
one has $\tilde{R}_{\rm m}^{+}=\tilde{R}_{\rm m}^{-}$ and the extrema vanishes. In this case, cloud can be unstable, and eventually collapse, leading to a bosenova. When $\tilde{\lambda}\gg 1$, we have $\epsilon_{\rm collapse}\approx(12I_{p}\alpha\tilde{\lambda}^{2})^{-1}$, which can be very small, indicating the cloud will collapse easily if the coupling is strong.
Figure 10: Numerical values of each angular momentum component of $I_{p}(q,\alpha)$. Since the integral is not significantly dependent on $\alpha$, we take $\alpha=0.1$ here as an example.
V Implications on Observations
According to Sec. II and Sec. IV, superradiance evolution may change dramatically in the presence of a second field. In particular, the superradiant efficiency is typically suppressed if the superradiant field couples to another field even with a very weak coupling strength. This effect has important implications on the observational signatures of superradiance, and should be considered in the dark particle searches based on superradiance. For instance, in Fig. 11, we show the spin of rapid rotating black holes after 10 Gyrs, assuming there are two interacting scalar fields with masses $\mu=10^{-12}\mathrm{eV}$ and $q=0.1$. Depending on the coupling strength, we find the mass-spin distribution can be quite different from the case of a single free field.
Figure 11: Mass-spin distribution of black holes after 10 Gyrs superradiant evolution. The red lines show the mass-spin distribution, assuming there are two interacting scalar fields of masses $\mu=10^{-12}\,\mathrm{eV}$ and $q=0.1$ and coupling with $\lambda=$ respectively, while the dashed gray line shows the distribution assuming a free field. The initial spin of black holes is $a_{0}=0.99$.
Even with a sufficient growth of the cloud, the existence of the other fields can change the evolution of the cloud, including its maximum mass, decay rate and lifetime, and hence leads to different gravitational wave signals comparing to the single field case. These effects should be taken into account in when searching for dark particles with black hole superradiance.
VI Conclusion and Discussion
In this work, we investigated black hole superradiance of interacting multiple fields. Taking a two-field model for demonstration, we discuss the possible interactions between the bound and unbound states of the two fields, and their impacts on the superradiant growth via a perturbative approach. We further analyze the superradiant growth of the most fastest mode in the presence of the interactions. We find that the superradiant mode typically experiences three stages: the exponential growth stage, the quasi-equilibrium stage and the power-law decay stage. We also find that the superradiance is typically suppressed if the superradiant field couples to another field even with a very weak coupling strength.
While we taking a two-field model for demonstration, inclusion of more fields should not alter the picture essentially. This is especially the case when there is a large hierarchy in the field mass such that one of the fields grows efficiently, and another field contributes most to the interaction corrections. Nevertheless, one could consider the superradiant growth of the other modes, for example, ${\bar{\psi}}_{322}$, ${\bar{\varphi}}_{211}$ or ${\bar{\varphi}}_{322}$ in our setup, which may play a role in the late evolution of ${\bar{\psi}}_{211}$ in certain parameter regions. Including such modes into the discussion is beyond the scope of this work, and will be studied in future work. It is also interesting to extend the discussion to the spin-1/2 and spin-1 fields, and to consider other possible interactions and their effects on superradiant growth.
It is possible that the dark sector consists of many species of interacting dark particles, rather than just one species of dark particle. As we demonstrated in this work that the existence of the other fields can have important impacts on the observational signatures of superradiance, the superradiance based constraints on dark particles derived from single-field analyses should be revised in the presence of interactions. The observational consequences of multi-field black hole superradiance, such as black hole spin distribution, gravitational wave emission from the clouds, and other potential signals of superradiant cloud, are also interesting questions to be studied in future.
ACKNOWLEDGMENTS
J. Z. is supported by the National Natural Science Foundation of China (NSFC) under Grants No. E414660101 and No. 12147103, and the Fundamental Research Funds for the Central Universities under Grants No. E4EQ6604X2 and No. E3ER6601A2. Y.-S. P. is supported by National Key Research and Development Program of China under Grant No.2021YFC2203004 and the NSFC under Grant No.12475064.
Appendix A Stability of quasi-equilibrium
In Sec. IV.1 and Sec. IV.2, we have discussed the asymptotical behavior near $a_{\rm crit}$ and the condition that the quasi-equilibrium becomes unstable. In this appendix, we will show more detailed derivation and generalize our discussion.
Consider a radiation channel with a $k$-power law, we have
$\dot{\epsilon}$ $=\gamma_{sr}(a)\epsilon-\gamma_{k}\epsilon^{k+1}$ (59)
$\dot{a}$ $=-\gamma_{sr}(a)\epsilon\,$,
where $k\geq 1,k\in\mathbb{Z}$ and $\gamma_{k}$ is a constant. Assuming the system is in equilibrium, we can substitute the equilibrium occupation number $\epsilon_{\rm eq}(a)=(\gamma_{sr}/\gamma_{k})^{1/k}$ into Eq. (59) and get
$\mathrm{d}\eta=-\frac{\mathrm{d}a}{\epsilon_{\rm eq}^{k+1}}\,$, (60)
where $\eta=\gamma_{k}\tilde{t}$ is the rescaled time. Integrating both sides of the equation and expand $a$ around $a_{\rm crit}$, we have
$\eta=k(a-a_{\rm crit})^{-\tfrac{1}{k}}\left(\frac{\alpha^{8}}{24\gamma_{k}}\frac{1+4\alpha^{2}}{1-4\alpha^{2}}\right)^{-\tfrac{1+k}{k}}\,$, (61)
and $\epsilon_{\rm eq}$ can be expanded by
$\epsilon_{\rm eq}=\left(\frac{a-a_{\rm crit}}{\gamma_{k}}\frac{\alpha^{8}}{24}\frac{1+4\alpha^{2}}{1-4\alpha^{2}}\right)^{\tfrac{1}{k}}\,.$ (62)
Hence the occupation number and black hole spin evolve as
$a-a_{\rm crit}$ $=\left(\frac{k}{\eta}\right)^{k}\left(\frac{24\gamma_{k}}{\alpha^{8}}\frac{1-4\alpha^{2}}{1+4\alpha^{2}}\right)^{k+1}$ (63)
$\epsilon_{\rm eq}$ $=\frac{24\gamma_{k}}{\alpha^{8}}\frac{1-4\alpha^{2}}{1+4\alpha^{2}}\frac{k}{\eta}\,.$
However, the quasi-equilibrium stage could be unstable. Considering a small deviation $\delta\epsilon$ from $\epsilon_{\rm eq}$, the evolution of $\delta\epsilon$ is given by
$\delta\epsilon^{\prime}$ $=\frac{\gamma_{sr}}{\gamma_{k}}(\epsilon_{\rm eq}+\delta\epsilon)-(\epsilon_{\rm eq}+\delta\epsilon)^{k+1}$ (64)
$=-k\epsilon_{\rm eq}^{k}\delta\epsilon\,$,
which means a typical timescale for restoring equilibrium is $\tau_{re}\sim(k\epsilon_{\rm eq}^{k})^{-1}$. On the other hand, $\epsilon_{\rm eq}$ varies on a time scale of
$\tau_{var}\sim\left|\frac{\epsilon_{\rm eq}}{\epsilon^{\prime}_{\rm eq}}\right|=\left|\frac{\epsilon_{\rm eq}}{\frac{{\rm d}\epsilon_{\rm eq}}{{\rm d}a}a^{\prime}}\right|=\left|\frac{\mathrm{d}\epsilon_{\rm eq}}{\mathrm{d}a}\epsilon_{\rm eq}^{k}\right|^{-1}\,.$ (65)
Therefore, we expect that equilibrium breaks down when $\tau_{re}\gtrsim\tau_{var}$, as the evolution of real occupation number $\epsilon$ can not catch up with the change of $\epsilon_{\rm eq}$. In case A, we have $k=1$, and $\tau_{re}<\tau_{var}$ is always satisfied when $a$ is close to $a_{\rm crit}$. For $k\geq 1$, there could be threshold in terms of black hole spin $a_{th}$, below which one has $\tau_{re}>\tau_{var}$, indicating the equilibrium is unstable. After the black hole spin drops below $a_{th}$, the power law decay of $\epsilon$ transforms from $t^{-1}$ to $t^{-1/k}$.